OpenStudy (anonymous):

simplify the square root. -64 will give medal

OpenStudy (mathmale):

Remember that the square root of -1 is i. Note that \[\sqrt{-64}\] can be re-written as \[\sqrt{-1}\sqrt{64}\]

OpenStudy (mathmale):

Hope this background info helps you to answer this question.